OpenStudy (anonymous):

Find the x value for point F such that DF and EF form a 1:3 ratio. E= (2,4) and D= (-3,-3) A) 0.75 B) −1 C) 2 D) −1.75

OpenStudy (anonymous):

 \[x = \frac{ mx2+nx1 }{ m+n } , y = \frac{ my2+ny1 }{ m+n }\]